Jordan Inc. manufactures Product B, incurring variable product costs of $15.00 per unit and fixed product costs of $70,000. Total Selling and Administrative Expenses for the year are expected to be $15,700. Jordan desires a profit equal to a 10% rate of return on assets, $785,000 of assets are devoted to producing Product B, and 100,000 units are expected to be produced and sold.

Compute the markup percentage, using the product cost concept.

Compute the normal selling price of Product B.
Answer #1

Answer 1:

Required Return:

Jordan desires a profit equal to a 10% rate of return on assets.

Value of assets that are devoted to producing Product B = $785,000

Required return = 785000 * 10% = $78,500

Product costs For 100,000 units:

Variable product costs = 100000 * $15 = $1,500,000

Fixed product costs = $70,000

Product cost = $1,570,000

Markup percentage:

Selling and Administrative Expenses = $15,700

Return/Profit required = $78,500

Markup percentage over product cost has to cover both 'Selling and Administrative Expenses' and 'profit required'

Markup percentage = (15700 + 78500) / 1570000 = 6%

The markup percentage, using the product cost concept = 6%

Answer 2:

Sale value for 100,000 units = Product costs + Selling and Administrative Expenses + Return required

= 1570000 + 15700 + 78500

= $1,664,200

Selling price of Product B = 1,664,200 / 100000 = $16.642

Selling price of Product B = $16.642 or $16.64